OpenStudy (camerondoherty):

How do you complete the square and find the vertex of: \[f(t)=−16t^2−32t+128\]

OpenStudy (wilder.monday):

Can you factor out a GCF ( greatest common factor)

OpenStudy (camerondoherty):

No, I've already factored it, I need to complete the square...

OpenStudy (camerondoherty):

Don't you have to first set the equation to zero and get rid of any constants?

OpenStudy (wilder.monday):

May I see all the work you did?

OpenStudy (wilder.monday):

First set it equal to y. Then, subtract 128 from both sides the left side should look like y - 128

OpenStudy (displayerror):

Divide the entire equation by the coefficient in front of the t^2 term: \[f(t) = t^2 + 2t - 8\] Move the 8 to the other side: \[t^2 + 2t = 8\] Divide the coefficient in front of the t term by 2, then square that result \[(\frac{2}{2})^2 = 1\] Add the squared result (1) to both sides of the equation \[t^2 + 2t + 1= 8+1\] Factor \[(t+1)^2= 9\] Solve for t \[t = \pm \sqrt{9} - 1\] \[t = -4, 2\] Find the corresponding y value for t=-4 and t=2

OpenStudy (wilder.monday):

Is this multiple choice? I just want to make sure what I'm doing goes along its the choices.

OpenStudy (acxbox22):

for the vertex you can find the the x value of the vertex by using x=-b/2a from the original equation then sub that x value into the equation to get the y value of the vertex

OpenStudy (camerondoherty):

To find the y value, do you sub the value for y and solve?

OpenStudy (camerondoherty):

*t

OpenStudy (displayerror):

Yeah, we found two values of t, so plug into f(t) to find your y value(s).

OpenStudy (camerondoherty):

Ok, so: \[f(t)=−16t^2−32t+128 \]\[f(t)=−16(-4)^2−32t+128 \]and\[f(t)=−16(2)^2−32t+128? \]

OpenStudy (camerondoherty):

Sheesh, Openstudy has a real character issue If you see diamonds, refresh so it goes back to normal cx

OpenStudy (acxbox22):

@DisplayError found the solutions now we have to find the vertex for vertex we can use x=-b/2a x= 32/(-16)(2) x=32/-32 x=-1 now sub -1 for x into the equation to find y or f(t)

OpenStudy (displayerror):

Oh wait, my bad. It was already assumed in completing the square that the equation was equal to 0 (completed the square to find the zeroes of the function), so we don't have to plug the x-values back into f(t)--they're 0 already.

OpenStudy (camerondoherty):

\[f(t)=−16t^2−32t+128 \] \[f(t)=−16(-1)^2−32t+128 \] \[f(t)=−16(1)−32t+128 \] \[f(t)=−16−32t+128 \] \[f(t)=−32t+112 \]

OpenStudy (acxbox22):

sorry @camerondoherty in this case x is t but it is the same thing plug in -1 for t

OpenStudy (camerondoherty):

I did

OpenStudy (acxbox22):

you forgot to plug it in for the term -32t

OpenStudy (camerondoherty):

Oh, lol, sorry cx Just noticed

OpenStudy (camerondoherty):

\[f(t) = -16 - 32(-1) + 128\] \[f(t) = -16+32 + 128\] \[f(t) = 16+ 128\] \[f(t)=144 \]

hartnn (hartnn):

let me give the steps to complete the square, see if you guys did the same thing, \(\Large x^2+bx+c \\ = \Large (x^2 + bx \color{blue}{+b^2/4}) + (c\color{blue}{-b^2/4}) \\ \Large = (x+(b/2))^2 +(c-b^2/4)\) have you guys done the similar thing ?

OpenStudy (acxbox22):

yes so we have the x and y values of the vertex now @camerondoherty where x=-1 and y/f(t) = 144 so the vertex is (-1,144)

OpenStudy (camerondoherty):

Ohhhh ok, Thank You c:

OpenStudy (acxbox22):

no problem :)